OpenStudy (anonymous):

WHAT VALUE OF C MAKES THE POLYNOMIAL A PERFECT SQUARE TRINOMIAL 4X^2-28X+C

OpenStudy (mathstudent55):

I don't understand the question. Can you check on the exact wording of the problem?

OpenStudy (mathstudent55):

Is the word "root" supposed to be there?

OpenStudy (anonymous):

I CHANGED IT

OpenStudy (mathstudent55):

Good. A perfect square trinomial is like the following two patterns: \((a + b)^2 = a^2 + 2ab + b^2 \) \((a - b)^2 = a^2 - 2ab + b^2 \)

OpenStudy (mathstudent55):

Since your middle term is -28x, we are following the second pattern.

OpenStudy (anonymous):

OKAY

OpenStudy (mathstudent55):

Notice that in the second pattern, the trianomial \(a^2 - 2ab + b^2 = (a - b)^2 \) In the trinomial you have \(a^2)\, but the squared binomial has only a, as in a + b.

OpenStudy (anonymous):

DONT WANT TO PRESSURE YOU BUT CAN YOU HURRY ITS BECAUSE I ONLY GET CURTAIN AMOUNT OF TIME FOR EACH PROBLEM

OpenStudy (mathstudent55):

We can start writing what the binomial has to be: \(4x^2 - 28x + c = (2x - ~?~)^2 \)

OpenStudy (mathstudent55):

We need to figure out what goes in the ?

OpenStudy (mathstudent55):

It has to be a number that when you square you get c.

OpenStudy (anonymous):

16

OpenStudy (anonymous):

OR 13

OpenStudy (mathstudent55):

Also, the middle term of the trinomial is -2ab, so -2ab = -28x, and a = 2x -2ab = -2(2x)b = -28x -4xb = -28x xb = 7x b = 7 Since b = 7, c = 49 \(4x^2 - 28x + 49 = (2x - 7)^2\) c = 49

OpenStudy (anonymous):

WOULD IT BE (2X-7)^2 OR (2X+7)(2X-7)

OpenStudy (mathstudent55):

We are not dealing with a difference of squares which is the product of a sum and a difference. This is the square of a binomial, so it is \( (2x - 7)^2 \)